OpenStudy (anonymous):

Hello Physics universe. I have a question about the kinematics equations. Could someone please explain how to derive them for constant acceleration

OpenStudy (anonymous):

dv/dt = a integrate v = a t + c ds/dt = v = at + c integrate s = (1/2)a t^2 + c t + e get the constants c and e from initial conditions usually they are initial position and initial velocity